A person goes shopping and spends 75% of his money. If he is now left with Rs. 600, find out how much he had in the beginning.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, 72% of 25 students like maths.

Hence, 72% of 25 = (72/100) × 25 = 18 students

Now, from 25 students, 18 students like maths

Thus, the number of students who do not like maths = 25 – 18 = 7 students

Thanks